find the volume of spheres whose surface area is 154 cm2
1
Expert's answer
2013-01-09T09:10:23-0500
Let R - radius of sphere then surface area=4*pi*R^2 inour situation

154=4pi*R^2
R^2=154/(4pi)=77/(2pi)
R=sqrt( 77/(2pi) )
Volume=4/3 *pi*R^3=4/3 *pi*(sqrt( 77/(2pi) ) )^3=4pi/3* 77/(2pi) *sqrt( 77/(2pi) )=154/3 *sqrt( 77/(2pi) )